# Dynamical Moiré Systems in Twisted Bilayer Optical Lattices

### v3

@@ Abstract
Figure 3: (a) Phase diagram with respect to $\rho_{\uparrow}$ and $J/U(g_{\sigma\sigma})$ with other parameters fixed at $g_{\uparrow\downarrow}/g_{\sigma\sigma}=0.9$, $\theta=36.87^{\circ}$, and $\rho_{\downarrow}=4/5$. (b) Phase diagram with respect to the twist angle $\theta$ and $J/U(g_{\sigma\sigma})$ with other parameters fixed at $g_{\uparrow\downarrow}/g_{\sigma\sigma}=0.9$, $\rho_{\uparrow}=1/10$, and $\rho_{\downarrow}=3/5$. (c) and (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dots and the blue squares in (a) and (b), respectively. See text for more details.
@@
In light of the tremendous experimental progress in ultracold atomic physics, the rich physics in this dynamical moiré system can be readily observed using current experimental setups Meng et al. (2023). Specifically, $J/U(g_{\sigma\sigma})$ can be tuned by adjusting the depth of optical lattice potential. The interlayer interaction $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)$ can be altered by either tuning the vertical displacement between the two spin-dependent lattices, or by utilizing the Feshbach resonance  [^2] The Feshbach resonance technique enables the direct tuning of $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ in the phase diagrams. For the interlayer interaction $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\equiv\frac{g_{\uparrow\downarrow}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\uparrow}}(\mathbf{r})\right|^{2}\left|w_{\mathbf{j}_{\downarrow}}(\mathbf{r})\right|^{2}$, displacing the spin-dependent lattices in out-of-plane direction reduces the overlap between Wannier functions $w_{{\uparrow/\downarrow}}(\mathbf{r})$, whose effect is physically equivalent to the change of $g_{\uparrow\downarrow}/g_{\sigma\sigma}$. The control over the twist angle can be realized by experimental techniques developed recently for the twisted bilayer optical lattices Meng et al. (2023). The exotic SF, MI phases, and various self-localization behavior can be directly observed via the quantum gas microscope Gross and Bakr (2021).
@@
Conclusions.—We propose a new type of moiré system in twisted bilayer optical lattices, namely, the dynamical moiré system that hosts a wide range of unique physical phenomena. It can not only accommodate moiré SF phase similar to the one found in conventional static bosonic moiré system, but also give rise to rich physics without static counterparts. These include “interlocked” MI, exotic SF and MI phases breaking moiré lattice symmetry, phases with one layer in SF and the other in MI, localization behavior similar to Bose glass and quasi-MBL without extrinsic disorder. Since both the dynamical moiré system itself and its experimental realization is even simpler than the conventional bosonic moiré system realized in experiments recently Meng et al. (2023), we believe our work will stimulate both further theoretical research on this type of dynamical moiré system, especially the possible intrinsic connection between the self-localization found here and quasi-MBL, generalization to fermionic systems, as well as experimental effort in searching for these exotic quantum phases.
@@ Acknowledgements.
Supplemental Material for “Dynamical Moiré Systems in Twisted Bilayer Optical Lattices”
I System and model
Motivated by the recent realization of the twisted bilayer optical lattices (Meng et al., 2023), we consider a closely related but simpler experimental set-up, namely, a twisted bilayer optical lattice system without the additional microwave field that has been used to realize the interlayer coupling in the recent experiments (Meng et al., 2023). More specifically, the system under study consists of two species of bosonic atoms that correspond to two hyperfine states (for instance, ${}^{87}\mathrm{Rb}$ atoms in hyperfine states $|F=1,m_{F}=1\rangle$ and $|F=2,m_{F}=0\rangle$ considered in Ref. (Meng et al., 2023)), with each species loaded in one layer of a state-dependent twisted bilayer square optical lattice, respectively [see Fig. 1(a) of the main text for a schematic illustration].
The physics of this system at low filling can be described by a Bose-Hubbard type model within the lowest band approximation. To derive the model Hamiltonian, we start with the following many-body Hamiltonian
$\hat{H}=$ $\sum_{\sigma=\uparrow,\downarrow}\int d^{3}\mathbf{r}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)\hat{\psi_{\sigma}}(\mathbf{r})$
$+\sum_{\sigma=\uparrow,\downarrow}\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r^{\prime}})\frac{V_{\sigma\sigma}(\mathbf{r}-\mathbf{r^{\prime}})}{2}\hat{\psi}_{\sigma}(\mathbf{r^{\prime}})\hat{\psi}_{\sigma}(\mathbf{r})$ (S1)
@@ I System and model
$V_{\uparrow\downarrow}(\mathbf{r}-\mathbf{r^{\prime}})$ $=\frac{4\pi\hbar^{2}a_{\uparrow\downarrow}}{\mu}\delta(\mathbf{r}-\mathbf{r^{\prime}}),$ (S2)
with $a_{\sigma\sigma}$ ($a_{\uparrow\downarrow}$) being the intra(inter)-species $s$-wave scattering length. The first term of the Hamiltonian (S1) is the single particle term, and the second term describes the intra-species interaction of atoms in the same lattice layer, and the third term describes the inter-species interaction between atoms of different lattice layer.
@@
$\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}=\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r}).$ (S3)
Therefore,
$\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})=\sum_{\mathbf{i}_{\sigma}}\langle\mathbf{i}_{\sigma}\mid\mathbf{r}\rangle\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}=\sum_{\mathbf{i}_{\sigma}}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger},$ (S4)
$\hat{\psi}_{\sigma}(\mathbf{r})=\sum_{\mathbf{i}_{\sigma}}\langle\mathbf{r}\mid\mathbf{i}_{\sigma}\rangle\hat{b}_{\mathbf{i}_{\sigma}}=\sum_{\mathbf{i}_{\sigma}}w_{\mathbf{i}_{\sigma}}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}},$ (S5)
where $\hat{b}_{\mathbf{i_{\sigma}}}^{\dagger}(\hat{b}_{\mathbf{i}_{\sigma}})$ is the creation (annihilation) operator at site $\mathbf{i}$ of the lattice layer $\sigma$ in the Wannier basis.
@@
$H_{\mathrm{s.p.},\sigma}=$ $\int d^{3}\mathbf{r}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)\hat{\psi_{\sigma}}(\mathbf{r})$
$=$ $\sum_{\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}}\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)w_{\mathbf{j}_{\sigma}}(\mathbf{r})\hat{b}_{\mathbf{j}_{\sigma}}$ (S6)
@@
$\epsilon_{\mathbf{i}_{\sigma}}\equiv\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})[\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})]w_{\mathbf{i}_{\sigma}}(\mathbf{r}),$ (S7)
@@
$J\equiv-\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})[\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})]w_{\mathbf{j}_{\sigma}}(\mathbf{r}).$ (S8)
@@
$H_{\mathrm{intra},\sigma}=$ $\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r^{\prime}})\frac{V_{\sigma\sigma}(\mathbf{r}-\mathbf{r^{\prime}})}{2}\hat{\psi}_{\sigma}(\mathbf{r^{\prime}})\hat{\psi}_{\sigma}(\mathbf{r})$
$=$ $\sum_{\mathbf{i}_{\sigma},\mathbf{j}_{\sigma},\mathbf{k}_{\sigma},\mathbf{l}_{\sigma}}\frac{1}{2}\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})w_{\mathbf{j}_{\sigma}}^{*}(\mathbf{r^{\prime}})\frac{4\pi\hbar^{2}a_{\sigma\sigma}}{\mu}\delta(\mathbf{r}-\mathbf{r^{\prime}})w_{\mathbf{k}_{\sigma}}(\mathbf{r^{\prime}})w_{\mathbf{l}_{\sigma}}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{j}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{k}_{\sigma}}\hat{b}_{\mathbf{l}_{\sigma}}$
$=$ $\sum_{\mathbf{i}_{\sigma}}\frac{U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}}{2}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}\hat{b}_{\mathbf{i}_{\sigma}}+\sum_{\left\langle\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}\right\rangle}\frac{U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}}{2}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}\hat{b}_{\mathbf{j}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{j}_{\sigma}}$ (S9)
@@
$U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}\equiv\frac{g_{\sigma\sigma}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\sigma}}(\mathbf{r})\right|^{4},$ (S10)
@@
$U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}\equiv\frac{g_{\sigma\sigma}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\sigma}}(\mathbf{r})\right|^{2}\left|w_{\mathbf{j}_{\sigma}}(\mathbf{r})\right|^{2}.$ (S11)
@@
$H_{\mathrm{inter}}=\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}},$ (S12)
@@
$U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\equiv\frac{g_{\uparrow\downarrow}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\uparrow}}(\mathbf{r})\right|^{2}\left|w_{\mathbf{j}_{\downarrow}}(\mathbf{r})\right|^{2},$ (S13)

### v4

@@ II System and model
Figure 2: Phase diagrams and typical configurations at the balanced and imbalanced fillings. (a) Phase diagram with respect to the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ and the hopping amplitude $J/U(g_{\sigma\sigma})$ at the balanced filling with $\rho_{\uparrow}=4/5$ and $\rho_{\downarrow}=4/5$ (twist angle is fixed at $\theta=36.87^{\circ}$). The magenta region of the phase diagram contains the BG phase. (b) Similar to (a), but at the imbalanced filling with $\rho_{\uparrow}=1/5$ and $\rho_{\downarrow}=4/5$. (c) and (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dots and the blue squares in (a) and (b), respectively.
III Results
III.1 Conventional moiré physics without interlayer tunneling
In conventional bosonic moiré systems, the interlayer tunneling causes layer hybridization of single particle orbitals, and thus particle filling on the hybridized orbitals results in phases with two layer sharing the same character. Indeed, as observed in recent experiments Meng et al. (2023), it can accommodate SF-SF and MI-MI (denoting SF/MI phase in $\uparrow$/$\downarrow$ / ↓ layer), with the particle density assuming the moiré lattice symmetry. In fact, these observations can be easily explained by treating the system as if the atoms on one layer experience an additional static potential from the other layer of the lattice Meng et al. (2023). Speculating along this line, we notice that although there is no interlayer tunneling in the dynamical moiré system studied here, atoms residing on one layer can nevertheless feel an additional potential provided by the atoms residing on the other layer via the interlayer interactions $\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}}$. This indicates the system may accommodate similar conventional moiré physics even without the interlayer tunneling.
In Fig. 2(a), we map out the phase diagram of the system in a balanced filling case with $\rho_{\uparrow}=\rho_{\downarrow}=4/5$ ($\rho_{\sigma}\equiv\langle\sum_{\mathbf{i}_{\sigma}}\hat{n}_{\mathbf{i}_{\sigma}}\rangle/N_{\sigma}^{\mathrm{lat}}$ with $N_{\sigma}^{\mathrm{lat}}$ being the number of lattice sites of the $\sigma$-layer within the cylinder trap). In the relatively large hopping regime, we can see that the system indeed support SF-SF phase (denoted as SF in the phase diagrams) similar to conventional static moiré systems, with both layers manifesting the SF order parameter $\phi_{\mathbf{i}_{\sigma}}\equiv\langle\hat{b}_{\mathbf{i}_{\sigma}}\rangle$ and density distribution $\langle\hat{n}_{\mathbf{i}_{\sigma}}\rangle$ assuming the moiré lattice symmetry [see Fig. 2(c1)]. While in the relatively small hopping regime, we notice that despite the filling factor on each layer being non-integer, increasing the inter-species interaction strength $g_{\uparrow\downarrow}$ can drive the system into the induced MI phase characterized by regions with integer-filled particles and vanishing SF order parameter in either layer [see Fig. 2(c4)], including MI-SF, MI-MI and SF-MI phases. This is in sharp contrast to the single-component Bose gases in optical lattices, where the MI can only exist at integer fillings.
The induced MI phase occurs because increased $g_{\uparrow\downarrow}$ enhances the effective potential experienced by atoms on one layer imposed by atoms on the other layer, causing localization. This scenario is reminiscent of conventional bosonic moiré systems, where increasing the interlayer tunneling makes one layer impose a stronger additional static lattice potential on the other layer, leading to stronger localization Meng et al. (2023). However, it’s important to recognize a key difference here: the effective potential is a dynamical one determined by the dynamical variables, and hence does not necessarily share the same spatial symmetry of the underlying optical lattice. For balanced filling case, indeed, we have not found MI phase with both layer assuming the moiré lattice symmetry. Instead, as shown in the Figs. 2(c2) and (c3), we find this dynamical moiré system can manifest rich discrete rotational symmetric MI and SF phases with no counterparts in the conventional bosonic moiré systems. However, away from the balanced filling, we indeed find the system can support MI phase with density distributions on both layers assuming the moiré lattice symmetry. For instance, as shown in the Fig. 2(d4), this kind of moiré MI phase can emerge in the induced MI regime of the system at an imbalanced filling $\rho_{\uparrow}=1/5,\rho_{\downarrow}=4/5$. In this case, $\uparrow$-particles occupy overlapping sites while $\downarrow$-particles occupy non-overlapping sites within each moiré period [see the Fig. 1(c)], minimizing interlayer repulsion and preserving lattice symmetry. If the filling factors deviate from $(\rho_{\uparrow}=1/5,\rho_{\downarrow}=4/5)$ for $\theta=36.87^{\circ}$, the remaining particles or vacant sites are likely to break the moiré lattice symmetry. Similarly, the system with other twist angle [e.g. the $(\theta=22.62^{\circ},\rho_{\uparrow}=2/5,\rho_{\downarrow}=3/5)$ case shown in the lower left plot of Fig. 1(d)] also hosts this kind of “interlocked” MI with moiré lattice symmetry. By contrast, for a SF phase with moiré lattice symmetry, the filling factor requirements are less restrictive because the number of atoms per site is not constrained to integer values, allowing particles to distribute themselves in accordance with the moiré lattice symmetry. We also notice that the interplay between interlayer repulsion and filling imbalance gives rise to the difference in the critical $J/U(g_{\sigma\sigma})$ shown in Figs. 2(a) and (b): in the imbalanced case ($\rho_{\uparrow}=1/5$), the larger interparticle distance on the $\uparrow$-layer requires a higher $J/U(g_{\sigma\sigma})$ for coherence and transition into the SF phase compared to the balanced case ($\rho_{\downarrow}=4/5$).
Although the dynamical moiré system can manifest SF or MI phase similar to the ones found in conventional bosonic moiré systems Meng et al. (2023), there are fundamental differences. Instead of being coupled by interlayer tunneling, the atoms on different layers are coupled by the interlayer density-density interactions, which preserve the independent U(1) symmetries in each layer. Consequently, the total particle number of each layer is conserved and can be different. These fundamental differences thus naturally indicate much richer physics with no counterparts in conventional bosonic moiré systems can emerge, as we shall discuss below.
III.2 Beyond the conventional moiré physics
Since the total particle number of each layer is conserved separately, the filling difference between the two layers in fact provides an extra tuning knob of new physics compared with the conventional bosonic moiré systems, as showcased in Fig. 2. In Fig. 3(a), we systematically investigate the influences of the filling difference by keeping the filling factor of the $\downarrow$-layer fixed at $\rho_{\downarrow}=4/5$ and changing the filling factor of the $\uparrow$-layer $\rho_{\uparrow}$ from $1/10$ to $4/5$. As we can see from the Fig. 3(c1), when the filling of the $\uparrow$-layer is small ($\rho_{\uparrow}=1/10$), the $\downarrow$-layer is in a SF phase and it imposes strong influences on atoms on the $\uparrow$-layer by localizing them and making them in induced MI phase despite $\rho_{\uparrow}=1/10$ being non-integer. As the $\rho_{\uparrow}$ gradually increases [see Figs. 3(c2) and (c3)], the $\uparrow$-layer also imposes an enhanced localization effects on the atoms on the $\downarrow$-layer by driving a SF-MI transition on this layer. We also observe that the average localization effect experienced by each atom in the $\uparrow$-layer decreases, making the $\uparrow$-layer transition into a SF phase at relatively high filling with $\rho_{\uparrow}=1/2$. This phenomenon is reminiscent of what happens in high-temperature superconductors, where doping can lead to significant changes in the system’s transport properties Keimer et al. (2015).
Similar to the conventional moiré system, the twist angle also has a significant impact here, as clearly shown in Fig. 3(b). For the case at the filling $(\rho_{\uparrow}=1/10,\rho_{\downarrow}=3/5)$, in the intermediate hopping regime, we see that at a relatively small twist angle $\theta=7.63^{\circ}$, the system is in the SF-SF phase. As the twist angle increases to $\theta=22.62^{\circ}$, the atoms in the $\uparrow$-layer are localized due to the interaction from the $\downarrow$-layer and enter the induced MI phase, forming the phase MI-SF [see Fig. 3(d2)]. Similar situation is also found at even larger twist angle, for instance, $\theta=36.87^{\circ}$ [see Fig. 3(d3)].
Interestingly, we notice that density distributions in the Figs. 3(d2) and (d3) assume rather random pattern. In these cases, the atoms on $\downarrow$-layer in fact provide a “disordered” effective potential to the atoms lying on the $\uparrow$-layer. Thus, one naturally expects that this “disordered” effective potential quenches hopping of the $\uparrow$-atoms and makes them localized, which can be interpreted as the interlayer interaction induced “self-localization” between the atoms on the two layers. The presence of the “disordered” effective potential is associated with phases with the spontaneous breaking of $C_{4}$ rotational symmetry of the system [see Fig. 3(c1) and Fig. 3(d)], indicating that there are four degenerate ground states (see the Appendix D for the degenerate ground states).
Figure 3: (a) Phase diagram with respect to $\rho_{\uparrow}$ and $J/U(g_{\sigma\sigma})$ with other parameters fixed at $g_{\uparrow\downarrow}/g_{\sigma\sigma}=0.9$, $\theta=36.87^{\circ}$, and $\rho_{\downarrow}=4/5$. (b) Phase diagram with respect to the twist angle $\theta$ and $J/U(g_{\sigma\sigma})$ with other parameters fixed at $g_{\uparrow\downarrow}/g_{\sigma\sigma}=0.9$, $\rho_{\uparrow}=1/10$, and $\rho_{\downarrow}=3/5$. The magenta regions of the phase diagrams contain the BG phase. (c) and (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dots and the blue squares in (a) and (b), respectively.
Among the states with irregular density pattern, such as the Fig. 3(d3) and the last plot of Fig. 1(d), there exist states with disconnected superfluid islands, exhibiting the characteristics of BG (see the Appendix E for exact identification). Thus, this system supports BG without extrinsic disorder (see the magenta region corresponding to the BG phase in the phase diagrams shown in Figs. 2 and 3), in contrast to previous studies with either (quasi)disorder or quasicrystalline lattices Fisher et al. (1989); Fallani et al. (2007); Pasienski et al. (2010); Söyler et al. (2011); Gadway et al. (2011); D’Errico et al. (2014); Meldgin et al. (2016); Johnstone et al. (2021a); Ciardi et al. (2023). BG phase in our system is characterized by the presence of isolated, non-percolated superfluid islands, and usually surrounded by MI regions. Therefore, in the phase diagrams, the BG phase is located within the broader “induced MI” region, except the filling-angle-match case [see the Fig. 2 (b)] whose density distributions with the moiré lattice symmetry would not provide an effective disordered potential. Consistent with previous studies, we have found many metastable configurations associated with the BG phase in our system (see the Appendix D for the metastable configurations). As the presence of BG at ground state usually implies many-body localization (MBL) at excited states Abanin et al. (2019), this system also probably hosts (quasi-)MBL without extrinsic disorder Schiulaz et al. (2015); Yao et al. (2016); Smith et al. (2017); Karpov et al. (2021). Note that here either the BG or the possible quasi-MBL are found at commensurate twist angles, i.e., the system has translational invariance in the bulk, in contrast to previously investigated localization at incommensurate angles Huang and Liu (2019); Paul et al. (2023); Madroñero et al. (2025) or the interacting 2D Aubry-André model with quasi-periodicity Szabó and Schneider (2020); Štrkalj et al. (2022). The possible quasi-MBL here may be explained by a mechanism similar to the exponentially slow relaxation predicted in certain translation-invariant systems Schiulaz et al. (2015); Yao et al. (2016). Further exploration of the relation between the localization found here with MBL and the generalization to incommensurate twist angles will be an intriguing subject for future study.
@@ III.2 Beyond the conventional moiré physics
In light of the tremendous experimental advancements in ultracold atomic physics, we believe that the rich physics in this dynamical moiré system can be readily observed using current experimental setups Meng et al. (2023). More specifically, the $J/U(g_{\sigma\sigma})$ can be tuned by adjusting the depth of optical lattice potential. The interlayer interaction $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)$ can be altered by either tuning the vertical displacement between the two spin-dependent lattices, or by utilizing the Feshbach resonance  [^2] The Feshbach resonance technique enables the direct tuning of $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ in the phase diagrams. For the interlayer interaction $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\equiv\frac{g_{\uparrow\downarrow}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\uparrow}}(\mathbf{r})\right|^{2}\left|w_{\mathbf{j}_{\downarrow}}(\mathbf{r})\right|^{2}$, displacing the spin-dependent lattices in out-of-plane direction reduces the overlap between Wannier functions $w_{{\uparrow/\downarrow}}(\mathbf{r})$, whose effect is physically equivalent to the change of $g_{\uparrow\downarrow}/g_{\sigma\sigma}$. The control over the twist angle can be realized by experimental techniques developed recently for the twisted bilayer optical lattices Meng et al. (2023). The exotic SF, MI phases, and various self-localization behavior can be directly observed via the quantum gas microscope Gross and Bakr (2021).
IV Conclusions
@@
We have proposed a dynamical moiré system induced purely by interlayer particle interaction in twisted bilayer optical lattices, which hosts a wide range of unique physical phenomena. both the dynamical moiré system itself and its experimental realization is even simpler than the conventional bosonic moiré system realized in experiments recently Meng et al. (2023), we believe our work will stimulate further theoretical research on this type of dynamical moiré system, especially the possible intrinsic connection between the self-localization found here and quasi-MBL, generalization to fermionic systems, as well as experimental effort in searching for these exotic quantum phases. In addition, while not present in our purely repulsive interlayer system, other exotic quantum phases like pair superfluid and dimerized phases could exist in bilayer systems with attractive interlayer interactions Macia et al. (2014); Cinti et al. (2017), warranting further study.
@@ Acknowledgements.
Appendix A Model Hamiltonian and Gutzwiller method
The low physics of this system can be described by a Bose-Hubbard type model within the lowest band approximation. To derive the model Hamiltonian, we start with the following many-body Hamiltonian
$\hat{H}=$ $\sum_{\sigma=\uparrow,\downarrow}\int d^{3}\mathbf{r}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)\hat{\psi_{\sigma}}(\mathbf{r})$ (2)
$+\sum_{\sigma=\uparrow,\downarrow}\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}\left[\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r^{\prime}})\frac{V_{\sigma\sigma}(\mathbf{r}-\mathbf{r^{\prime}})}{2}\right.$
$\left.\hat{\psi}_{\sigma}(\mathbf{r^{\prime}})\hat{\psi}_{\sigma}(\mathbf{r})\right]$
@@ Appendix A Model Hamiltonian and Gutzwiller method
$V_{\uparrow\downarrow}(\mathbf{r}-\mathbf{r^{\prime}})$ $=\frac{4\pi\hbar^{2}a_{\uparrow\downarrow}}{\mu}\delta(\mathbf{r}-\mathbf{r^{\prime}}),$ (3)
with $a_{\sigma\sigma}$ ($a_{\uparrow\downarrow}$) being the intra(inter)-species $s$-wave scattering length. The first term of the Hamiltonian (2) is the single particle term, and the second term describes the intra-species interaction of atoms in the same lattice layer, and the third term describes the inter-species interaction between atoms of different lattice layer.
@@
$\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}=\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r}).$ (4)
@@
$H_{\mathrm{s.p.},\sigma}=$ $\int d^{3}\mathbf{r}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)\hat{\psi_{\sigma}}(\mathbf{r})$ (5)
$=$ $\sum_{\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}}\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)w_{\mathbf{j}_{\sigma}}(\mathbf{r})\hat{b}_{\mathbf{j}_{\sigma}}$
@@
$\epsilon_{\mathbf{i}_{\sigma}}\equiv\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})[\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})]w_{\mathbf{i}_{\sigma}}(\mathbf{r}),$ (6)
@@
$J\equiv-\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})[\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})]w_{\mathbf{j}_{\sigma}}(\mathbf{r}).$ (7)
@@
$H_{\mathrm{intra},\sigma}=$ $\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r^{\prime}})\frac{V_{\sigma\sigma}(\mathbf{r}-\mathbf{r^{\prime}})}{2}\hat{\psi}_{\sigma}(\mathbf{r^{\prime}})\hat{\psi}_{\sigma}(\mathbf{r})$
$=$ $\sum_{\mathbf{i}_{\sigma},\mathbf{j}_{\sigma},\mathbf{k}_{\sigma},\mathbf{l}_{\sigma}}\frac{1}{2}\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}\left[w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})w_{\mathbf{j}_{\sigma}}^{*}(\mathbf{r^{\prime}})\frac{4\pi\hbar^{2}a_{\sigma\sigma}}{\mu}\right.$
$\left.\delta(\mathbf{r}-\mathbf{r^{\prime}})w_{\mathbf{k}_{\sigma}}(\mathbf{r^{\prime}})w_{\mathbf{l}_{\sigma}}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{j}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{k}_{\sigma}}\hat{b}_{\mathbf{l}_{\sigma}}\right]$ (8)
$=$ $\sum_{\mathbf{i}_{\sigma}}\frac{U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}}{2}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}\hat{b}_{\mathbf{i}_{\sigma}}+\sum_{\left\langle\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}\right\rangle}\frac{U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}}{2}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}\hat{b}_{\mathbf{j}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{j}_{\sigma}}$
@@
$U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}\equiv\frac{g_{\sigma\sigma}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\sigma}}(\mathbf{r})\right|^{4},$ (9)
@@
$U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}\equiv\frac{g_{\sigma\sigma}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\sigma}}(\mathbf{r})\right|^{2}\left|w_{\mathbf{j}_{\sigma}}(\mathbf{r})\right|^{2}.$ (10)
@@
$H_{\mathrm{inter}}=\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}},$ (11)
@@
$U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\equiv\frac{g_{\uparrow\downarrow}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\uparrow}}(\mathbf{r})\right|^{2}\left|w_{\mathbf{j}_{\downarrow}}(\mathbf{r})\right|^{2},$ (12)